OpenStudy (anonymous):

what is the solution of the following system of equation? 2x-6=2y 3-2x=y

OpenStudy (anonymous):

x=2 and y=-1

OpenStudy (anonymous):

plug in 3-2x for y for the equation 2x-6

OpenStudy (anonymous):

right, substitute: you could do as bradley said: y = (3 - 2x) --> (2x - 6) = 2*(3 - 2x) --> (2x-6) = (6 - 4x)

OpenStudy (anonymous):

so 2x + 4x -6 = 6, therefore 6x - 6 = 6, therefore x = 12/6 = 2, x = 2, now use the fact that 3-2x=y --> y now is 3-2*(2) = 3-4 = -1 --> (x,y) = ( 2, -1 )

OpenStudy (anonymous):

ok. i am going to see if i can understand it so i can work on my next one

OpenStudy (anonymous):

so, just tell me how i sovled it?

OpenStudy (anonymous):

so you get the general concept

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

ok, what is confusing now?

OpenStudy (anonymous):

i guess the way it is written

OpenStudy (anonymous):

ok, you have two equations and two variables

OpenStudy (anonymous):

toussainr might be able to explain it differently...

OpenStudy (anonymous):

ok, well then: the goal is find the values for each variable that satisfy both equations

OpenStudy (anonymous):

My browser just froze, sorry

OpenStudy (anonymous):

dpflan, please go ahead!!!

OpenStudy (anonymous):

okay, I'm going to give it try, hoping I'm helpful

OpenStudy (anonymous):

you have to equations. we are going to proceed in steps to make things easier!!!

OpenStudy (anonymous):

1. You move terms in 'x' and 'y' to the left hand side (LHS) and the constants to the right hand side (RHS)

OpenStudy (anonymous):

So we have: 2x-2y = 6

OpenStudy (anonymous):

AND

OpenStudy (anonymous):

-2x-y = -3

OpenStudy (anonymous):

Let's move on with step two

OpenStudy (anonymous):

2. From equation "-2x-y = -3" Let's derive a value for 'y'

OpenStudy (anonymous):

We have -y = -3+2x OR y = 3-2x (we multiply by -1 the LHS and the RHS)

OpenStudy (anonymous):

This value: "y = 3-2x" we are going to replace it in the first equation : 2x-2y = 6

OpenStudy (anonymous):

We have now:

OpenStudy (anonymous):

2x - 2 (3-2x) = 6

OpenStudy (anonymous):

Let's multiply,. we have: 2x - 6 + 4x = 6

OpenStudy (anonymous):

Let's perform the required addition: 6x-6 = 6

OpenStudy (anonymous):

let's divide by 6 the RHS and the LHS, we have now : "x - 1 = 1"

OpenStudy (anonymous):

Form here we can easily find the value of x :" x = 2"

OpenStudy (anonymous):

Recall that y = 3-2x So, we replace: y = 3 - 2 * 2 ( * stands here for multiplication ) And finally y = 3 -4 y = -1

OpenStudy (anonymous):

THAT"S THE BEST I CAN DO!

OpenStudy (anonymous):

thank you. I understand it more now. :)

OpenStudy (anonymous):

Okay!!! Just to quote John Von Neuman : "You don't understand maths, you get accustomed to it" So do a many exercises as you can! That's the key to successful mathematics!